# On $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ and Frege Systems

### v4

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The techniques used in this paper are lazy-diagonalization and the novel new technique developed in author’s recent work [Lin21]. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there are language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ and other which is in the area of proof complexity implied by our main outcome. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomial bounded.

### v5

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The techniques used in this paper are lazy-diagonalization and the novel new technique developed in author’s recent work [Lin21]. Since there exists oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mysterious behind it and showing that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is not enumerable. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there are language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences implied by our main outcome such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ and other which is in the area of proof complexity. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomial bounded.
@@ 9 Conclusions
We have broken the so-called Relativization Barrier by proving that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is in fact not enumerable, thus showing that the ordinary diagonalization techniques will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.